What is the total charge output of the circuit below? 6F 3F CHE 12 F 6F 12 volt


1
Expert's answer
2021-07-13T10:18:49-0400

Gives


"C_1=6F,C_2=3F,C_3=12F,C_4=6F,V=12v"

all capacitor are connected series combination

"\frac{1}{C_{net}}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}"

"\frac{1}{C_{net}}=\frac{1}{6}+\frac{1}{3}+\frac{1}{12}+\frac{1}{6}"

"\frac{1}{C_{net}}=\frac{2+4+1+2}{12}"

"C_{net}=\frac{4}{3}"



Charge

"Q=CV"

Put value

"Q=\frac{4}{3}\times12=16C"